# A neural network-based approach to hybrid systems identification for control

### v1

@@ 6.1 Comparison with PARC
We first contrast the performance achieved by the proposed approach with that obtained by PARC, a state-of-the-art PWA regression method that has been configured to perform up to 25 iterations and consider 20 clusters. To take into account the nonconvex nature of both learning problems, each test has been repeated for 10 different initial conditions of the optimizer.
For reference, solving the training problem in (13) required less than 3 minutes on a laptop equipped with a intel core i7-1165G7 processor.
@@ 6.1.1 Piecewise system benchmarks
  • The problem adopted in [11, §IV.A], which we simulate from a random initial condition using an excitation signal with the features described in [11]. We will refer to this system, which has also been used as benchmark in [13, 53], as $\Sigma_{\mathrm{B-PWA}}$;
  • The queuing network described in [28, Ex. 5], in which we assume the second and third components of the vector $\mu$ are each restricted to the interval $[10,100]$, while its first component remains set to $30$ and the vector $s$ set to $[1000,11,11]$. The system is excited using a random white sequence drawn from a uniform distribution defined on the range above. The resulting ODEs are integrated using the LSODA suite [39] via Scipy. The sampling time is set to 5 seconds. We will refer to this system as $\Sigma_{\mathrm{QN}}$.
The numerical results for the examples above are reported in Table 1, which clearly show how our methodology works well, achieving performance comparable with PARC and, in turn, with the current state-of-the-art.
median BFR best BFR median RMS best RMS
$\Sigma_{\mathrm{PWA}}$ 0.952 (0.936) 0.959 (0.943) 0.0484 (0.0641) 0.0422 (0.057)
$\Sigma_{\mathrm{B-PWA}}$ 0.88 (0.852) 0.932 (0.855) 0.084 (0.109) 0.05 (0.08)
$\Sigma_{\mathrm{QN}}$ 0.838 (0.81) 0.851 (0.829) 0.192 (0.224) 0.175 (0.204)
Table 1: Numerical results obtained by our method and PARC (within parentheses) on PWA benchmarks.
@@ 6.1.2 Linear parameter-varying and nonlinear benchmarks
  • The linear parameter-varying dynamics in [11, §IV.B], excited according to the related discussion in that paper. We will refer to this system, which has also been used as benchmark in [13, 53], as $\Sigma_{\mathrm{B-LPV}}$;
@@
The numerical results for the examples above are reported in Table 2, where we can see that, on the $\Sigma_{\mathrm{B-LPV}}$ benchmark, our method greatly outperforms PARC. In addition, we note that the performance obtained on $\Sigma_{\mathrm{Tank}}$ are comparable with that in [51], with our methodology characterized by far fewer parameters and requiring much shorter training time than the NN in [51]. Finally, the results on $\Sigma_{2}$ are also very competitive w.r.t. those in [51], especially in view of the larger dataset employed in the latter paper (i.e., 20000 against 5000 samples).
median BFR best BFR median RMS best RMS
$\Sigma_{\mathrm{Tank}}$ 0.901 (0.9) 0.931 (0.91) 0.107 (0.109) 0.07 (0.09)
$\Sigma_{\mathrm{B-PLV}}$ 0.696 (0.511) 0.731 (0.528) 0.286 (0.46) 0.254 (0.44)
$\Sigma_{2}$ 0.924 (0.921) 0.947 (0.931) 0.061 (0.063) 0.042 (0.055)
Table 2: Numerical results obtained by our method and PARC (within parentheses) on the nonlinear and LPV benchmarks.
@@ 6.2 Sensitivity analysis – parameters $n_{r}^{\alpha}$, $n_{r}^{\beta}$
We now investigate the sensitivity of our NN-based approach w.r.t. the main parameters characterizing our technique, i.e., $n_{r}^{\alpha}$ and $n_{r}^{\beta}$. Although the diagonal Hessian matrices $Q_{\alpha}$, $Q_{\beta}$ represent further possible hyperparameters to tune, we note that according to both the discussion in [37, 38] and our own observation, the performance of our method is not particularly affected by their values. We therefore omit the related sensitivity analysis for these terms.
We report in Table 3 the median, worst and best BFR achieved on the $\Sigma_{2}$ and $\Sigma_{\mathrm{B-PWA}}$ for different values of $n_{r}^{\alpha}=n_{r}^{\beta}$. In general we can observe that, while a higher $n_{r}^{\alpha}$ is beneficial to improve the best-case scenario performance, it also makes the training problem more difficult. This is clearly shown by the obtained median BFR, especially when dealing with $\Sigma_{2}$. From our numerical experience, the difficulties of training models featuring higher values of $n_{r}^{\alpha}$ (and therefore the capability of representing richer continuous PWA mappings) can be often overcome by making use of larger training dataset. This behaviour is illustrated in Fig. 3, where we report a scatter plot of the relationship between the logarithm of the ratio between available samples and $n_{r}^{\alpha}$, and the median BFR achieved on $\Sigma_{\mathrm{B-PWA}}$. These findings also suggest that in practical settings in which the training data are fixed in advance, the problem of choosing appropriate values for $n_{r}^{\alpha}$ and $n_{r}^{\beta}$ may become central. To this end, the application of techniques such as group LASSO [25] or the one in [12] is currently under investigation.
Figure 3: Scatter plot of the logarithm of the ratio between the number of available samples (on the $x-$axis) and the median BFR (on the $y-$axis) when learning models using our proposed technique on $\Sigma_{\mathrm{B-PWA}}$.
$\Sigma_{\mathrm{2}}$ $\Sigma_{\mathrm{B-PWA}}$
median best median best
$n_{r}^{\alpha}=2$ 0.915 0.943 0.863 0.863
$n_{r}^{\alpha}=3$ 0.897 0.925 0.889 0.913
$n_{r}^{\alpha}=5$ 0.904 0.932 0.915 0.924
$n_{r}^{\alpha}=7$ 0.883 0.932 0.924 0.947
$n_{r}^{\alpha}=11$ 0.845 0.936 0.939 0.956
$n_{r}^{\alpha}=17$ 0.875 0.910 0.946 0.952
Table 3: Results obtained by our method in learning the behavior of $\Sigma_{\mathrm{2}}$ and $\Sigma_{\mathrm{B-PWA}}$ for different values of $n_{r}^{\alpha}$.
@@ 6.3 Performance for predictive control
Nevertheless, the reader might wonder if the proposed NN model can be used in an OCP scheme in conjunction with a sensitivity-based NLP solver. This is especially interesting for cases in which one is interested in working with a system as those introduced in §6.1.2 that do not exhibit behaviours involving logic and dynamic at the same time. To this end, we design a simple OCP for $\Sigma_{2}$, mimicking the setup of [51], with prediction horizon equal $T=7$. We solve, at each step, the following optimization problem:
@@
We then compare the behaviour of a closed-loop trajectory achieved when using SLSQP solver with the one obtained with the global optimization solver DIRECT_L [26]. The results are shown in Fig. 4 where, although the selected input sequence is slightly different, the closed-loop trajectories achieved by the two controllers are essentially indistinguishable. This suggests that derivative-based solver may also be capable of exploiting the proposed NN-based models. We note that on our reference hardware, the whole closed-loop simulation with the SLSQP-based controller required approximately one millisecond per time-step. Finally, we note that this test highlights the robustness of the surrogate NN model w.r.t. the covariate shift phenomena [59, Ch. 33.2.1].
@@
7 Conclusion
We have proposed a NN-based methodology to system identification for control that only requires the training of a NN via standard tools as simple identification step, and yields a hybrid system model suited for optimal control design. Specifically, we have employed a NN with specific, yet simple, architecture, which turns out to be end-to-end trainable and produces a hybrid system with PWA dynamics from available data. Extensive numerical simulations have illustrated that, as a NN-based identification procedure, our technique has very similar performance compared to the state-of-the-art of hybrid system identification methods. At the same time, we have shown that, under a careful choice of some weights of the NN, the resulting hybrid dynamics can be controlled (locally) optimally by just solving the KKT system associated to the underlying finite horizon OCP. This is computationally advantageous compared to traditional approaches to optimal control of hybrid systems, which usually require mixed-integer programming.

### v2

@@ 6.1 Comparison with PARC
We first contrast the performance achieved by our method with that obtained by piecewise-affine regression and classification (PARC), a state-of-the-art PWA regression method that has been configured to perform up to 25 iterations and consider 20 clusters. To take into account the nonconvex nature of both learning problems, each test has been repeated for 10 different initial conditions of the optimizer.
@@ 6.1.1 Piecewise system benchmarks
  • The problem adopted in [10, §IV.A], which we simulate from a random initial condition using an excitation signal with the features described in [10]. We will refer to this system, which has also been used as benchmark in [11, 47], as $\Sigma_{\mathrm{B-PWA}}$;
  • The queuing network described in [25, Ex. 5], in which we assume the second and third components of the vector $\mu$ are each restricted to the interval $[10,100]$, while its first component remains set to $30$ and the vector $s$ set to $[1000,11,11]$. The system is excited using a random white sequence drawn from a uniform distribution defined on the range above. The resulting ODEs are integrated using the LSODA suite [34] via Scipy. The sampling time is set to 5 seconds. We will refer to this system as $\Sigma_{\mathrm{QN}}$.
The numerical results for the examples above are reported in Table 1, which clearly show how our methodology works well, achieving performance comparable with PARC and, in turn, with the current state-of-the-art. From our numerical experience we note that solving the training problem in (13) on, e.g., the $\Sigma_{2}$ benchmark, required around $15$[s], while PARC approximately $30$[s].
median BFR best BFR variance BFR
$\Sigma_{\mathrm{PWA}}$ 0.952 (0.937) 0.959 (0.944) $2.88\cdot 10^{-4}$ ($2.66\cdot 10^{-5}$)
$\Sigma_{\mathrm{B-PWA}}$ 0.883 (0.855) 0.932 (0.891) $7.28\cdot 10^{-3}$ ($7.48\cdot 10^{-4}$)
$\Sigma_{\mathrm{QN}}$ 0.838 (0.81) 0.851 (0.829) $2.20\cdot 10^{-4}$ ($4.41\cdot 10^{-4}$)
median RMS best RMS variance RMS
$\Sigma_{\mathrm{PWA}}$ 0.048 (0.064) 0.042 (0.057) $2.95\cdot 10^{-4}$ ($3.12\cdot 10^{-5}$)
$\Sigma_{\mathrm{B-PWA}}$ 0.084 (0.109) 0.050 (0.082) $4.20\cdot 10^{-3}$ ($4.58\cdot 10^{-4}$)
$\Sigma_{\mathrm{QN}}$ 0.193 (0.224) 0.175 (0.205) $3.40\cdot 10^{-4}$ ($5.90\cdot 10^{-4}$)
Table 1: BFR and RMS obtained by our method and PARC (within parentheses) on PWA benchmarks.
@@ 6.1.2 Linear parameter-varying and nonlinear benchmarks
  • The linear parameter-varying dynamics in [10, §IV.B], excited according to the related discussion in that paper. We will refer to this system, which has also been used as benchmark in [11, 47], as $\Sigma_{\mathrm{B-LPV}}$;
@@
Table 2 reports the numerical values obtained for these examples, where we can see that, on the $\Sigma_{\mathrm{B-LPV}}$ benchmark, our method greatly outperforms PARC. On the reference hardware, for $\Sigma_{\mathrm{tank}}$ PARC takes $25$[s], while our training procedure around $40$[s]. On this benchmark, while the identification performance of our approach is comparable to that of [45], our training time is significantly lower than that reported in [45] (around $20$ minutes). This is mostly related to the fact that our method needs to learn a few hundreds of coefficients only. Finally, the results on $\Sigma_{2}$ are also very competitive w.r.t. those in [45], especially in view of the larger dataset employed in the latter paper (i.e., 20000 against 5000 samples).
median BFR best BFR variance BFR
$\Sigma_{\mathrm{Tank}}$ 0.901 (0.900) 0.932 (0.909) $4\cdot 10^{-4}$ ($1.64\cdot 10^{-3}$)
$\Sigma_{\mathrm{B-PLV}}$ 0.696 (0.511) 0.731 (0.528) $4.47\cdot 10^{-4}$ ($1.62\cdot 10^{-4}$)
$\Sigma_{2}$ 0.924 (0.921) 0.947 (0.931) $1.69\cdot 10^{-4}$ ($3.47\cdot 10^{-5}$)
median RMS best RMS variance RMS
$\Sigma_{\mathrm{Tank}}$ 0.108 (0.109) 0.075 (0.09) $4.76\cdot 10^{-4}$ ($1.95\cdot 10^{-3}$)
$\Sigma_{\mathrm{B-PLV}}$ 0.287 (0.46) 0.255 (0.44) $3.96\cdot 10^{-4}$ ($1.30\cdot 10^{-4}$)
$\Sigma_{2}$ 0.061 (0.063) 0.042 (0.056) $1.09\cdot 10^{-4}$ ($1.96\cdot 10^{-5}$)
Table 2: BFR and RMS obtained by our method and PARC (within parentheses) on the nonlinear and LPV benchmarks.
@@ 6.2 Sensitivity analysis – parameters $n_{r}^{\alpha}$, $n_{r}^{\beta}$
We now investigate the sensitivity of our NN-based approach w.r.t. the main parameters characterizing our technique, i.e., $n_{r}^{\alpha}$ and $n_{r}^{\beta}$. Note that also $Q_{\alpha}$, $Q_{\beta}$ represent further possible hyperparameters to tune, however, we note that according to both the discussion in [32, 33] and our own observation, the performance of our method is not particularly affected by their values. We therefore omit the related sensitivity analysis for these terms.
We report in Table 3 the median, worst and best BFR achieved on the $\Sigma_{2}$ and $\Sigma_{\mathrm{B-PWA}}$ for different values of $n_{r}^{\alpha}=n_{r}^{\beta}$. In general we can observe that, while a higher $n_{r}^{\alpha}$ is beneficial to improve the best-case scenario performance, it also makes the training problem more difficult. This is clearly shown by the obtained median BFR, especially when dealing with $\Sigma_{2}$. From our numerical experience, the difficulties of training models featuring higher values of $n_{r}^{\alpha}$ (and therefore the capability of representing richer continuous PWA mappings) can be often overcome by making use of larger dataset. These results also stress that, in case the date are fixed in advance, choosing appropriate values for $n_{r}^{\alpha}$ and $n_{r}^{\beta}$ is key.
$\Sigma_{\mathrm{B-PWA}}$ $\Sigma_{\mathrm{2}}$
median best variance median best variance
$n_{r}^{\alpha}=2$ 0.915 0.943 $1.06\cdot 10^{-3}$ 0.863 0.863 $1.21\cdot 10^{-3}$
$n_{r}^{\alpha}=3$ 0.897 0.925 $3.51\cdot 10^{-4}$ 0.889 0.913 $8.92\cdot 10^{-5}$
$n_{r}^{\alpha}=5$ 0.904 0.932 $3.23\cdot 10^{-3}$ 0.915 0.924 $7.05\cdot 10^{-5}$
$n_{r}^{\alpha}=7$ 0.883 0.932 $7.28\cdot 10^{-3}$ 0.924 0.947 $1.69\cdot 10^{-4}$
$n_{r}^{\alpha}=11$ 0.845 0.936 $5.99\cdot 10^{-3}$ 0.939 0.956 $2.85\cdot 10^{-4}$
$n_{r}^{\alpha}=17$ 0.875 0.910 $5.87\cdot 10^{-3}$ 0.946 0.952 $2.49\cdot 10^{-4}$
Table 3: BFR obtained by our method in learning the behavior of $\Sigma_{\mathrm{2}}$ and $\Sigma_{\mathrm{B-PWA}}$ for different values of $n_{r}^{\alpha}$.
@@ 6.3 Performance for predictive control
Nevertheless, one may wonder how the adopted NN model can be used within an OCP in conjunction with a sensitivity-based NLP solver. This is relevant for systems as those introduced in §6.1.2, which do not exhibit behaviors involving logic and dynamic at the same time. We then design a simple OCP for $\Sigma_{2}$, mimicking the setup of [45], with prediction horizon equal $T=7$. We solve, at each step, the following optimization problem:
@@
We then compare the behaviour of a closed-loop trajectory achieved when using SLSQP solver with the one obtained with the global optimization solver DIRECT_L [23]. The results are shown in Fig. 3 where, although the selected input sequence is slightly different, the closed-loop trajectories achieved by the two controllers are almost indistinguishable, thereby showing that derivative-based solver may exploit the adopted NN-based models.
We note that on our reference hardware, the whole closed-loop simulation with the SLSQP-based controller required approximately one millisecond per time-step. Overall, the SLSQP-based approach requires around $4$[s], in contrast to the $25$[s] needed by that based on DIRECT_L. While we opted for “off-the-shelf” solvers, adopting dedicated methodologies such as, e.g., [48, 20], would yield even better computational performance.
@@
7 Conclusion and outlook
We have proposed a NN-based methodology to system identification for control that only requires the training of a NN via standard tools as simple identification step, and yields a hybrid system model suited for optimal control design. Specifically, we have employed a NN with specific, yet simple, architecture, which turns out to be end-to-end trainable and produces a hybrid system with PWA dynamics from available data. Extensive numerical simulations have illustrated that, as a NN-based identification procedure, our technique has very similar performance compared to the state-of-the-art of hybrid system identification methods. By relying on available results [33], we have also shown that, under a careful choice of some weights of the NN, the resulting hybrid dynamics can be controlled locally optimally by just solving the KKT system associated to the underlying finite horizon OCP. This is computationally advantageous compared to traditional approaches to optimal control of hybrid systems, usually requiring mixed-integer optimization.
Future work will concentrate on the integration of OptNet layers together with standard explicit layers characterized by different activation functions. In view of the requirement in Corollary 1, it is also key to investigate the practical impact the various strategies to impose the conditions in (10) have on the resulting NN model.